18. Find the weight density (in lb / in3) of a block 5.00 in by 7.50 in x 9.00 in. that weighs 750. lb.
Answer #1

Volume = lbh = 5*7.5*9 = 337.5 in^3

Density = mass/Volume = 750/337.5 = 2.22 lb/in^3